# Data-driven optimization of pixelated CdZnTe spectrometers for uranium enrichment assay

### v1

@@ I Introduction
The H3D M400 gamma spectrometer [1] is being adopted by the International Atomic Energy Agency (IAEA) as its primary in-field uranium enrichment non-destructive assay (NDA) technology [2, 3, 4]. The M400 features four CdZnTe (CZT) crystals each pixelated to an $11\times 11$ grid, and offers medium resolution gamma spectroscopy at room temperature in a compact form factor. In previous work, we showed that the spatial variations in detector performance within the CZT crystal volumes could be exploited to optimize spectroscopic performance metrics [5]. In particular, a balance can be found between identifying and rejecting poorly-performing detector regions and the associated loss of efficiency. Because of the large number of possible voxel combinations (${\sim}10^{7285}$), computing the globally optimum set of voxels to include is infeasible, and approximate, data-driven methods were developed. Several of the example optimizations therein focused on minimizing the relative uncertainty in a peak fit amplitude parameter as proofs of concept in lieu of more advanced spectral performance metrics.
In this work, we have integrated our spectroscopic optimization software spectre-ml with GEM [6], the actual software used by the IAEA to conduct uranium NDA measurements. We show that the relative uncertainty on the GEM enrichment calculation can directly be used as a spectre-ml optimization target, i.e., that spectre-ml can improve the statistical confidence of the IAEA’s in-field enrichment measurements without increasing measurement times.
@@ II-A Integration of the spectre-ml and pyGEM codes
The spectre-ml [5, 7] code is a Python package for optimizing the performance of many-channel gamma spectrometers by intelligently rejecting data from poorly-performing detector channels (i.e., spatial regions). It uses non-negative matrix factorization (NMF) [8, 9] and unsupervised clustering algorithms to learn groups of voxels with similar performance, compute performance metrics across various voxel cluster combinations, and then sweep over hyperparameters such as clustering algorithm, number of clusters, number of NMF components, etc., to find the best overall set of voxels to include. Previous versions of spectre-ml were restricted to single Doniach [10] or Gaussian peak fit workflows using the becquerel library for peak fits. In this work we have generalized the spectre-ml software to abstract out the dependence on becquerel and enable arbitrary user-defined spectrum analyses, in particular those based on GEM. The pyGEM [11] code is a Python interface to the General Enrichment Measurements (GEM) code [6] used by the IAEA for uranium enrichment NDA. It fits measured spectra in the ${\sim}120$$270$ keV ROI using a set of known U-235 emission lines, forward scatter, and high-energy downscatter profiles, and has recently been updated to handle the asymmetric peak shapes of CZT. In contrast to our Doniach fit becquerel workflow, pyGEM uses a triple-Gaussian model for each peak in CZT—one primary peak, one low-energy tail, and one high-energy tail. The user provides pyGEM a spectrum from a sample of known U-235 enrichment, and it will then compute a linear calibration between the fit net count rate in the $185.7$ keV peak and the U-235 atom abundance that can be used to compute the enrichment from an unknown sample spectrum. It is also possible to incorporate correction factors for certain changes between the calibration and sample spectra such as changes in the sample matrix (e.g., U metal vs. U3O8), Al and steel wall thicknesses, etc.
@@
To integrate the two codes, some spectre-ml updates were required to handle both the becquerel and pyGEM workflows. As shown in the pseudocode in the Appendix, spectre-ml now provides the abstract base classes SpectrumAnalyzer and RankingMetric. The user must provide concrete subclasses implementing SpectrumAnalyzer.analyze and RankingMetric.calc, where the latter method takes a spectrum analyzed (e.g., fit) by the former and extracts a single float metric such as the relative uncertainty in a fit parameter (becquerel) or in an enrichment (pyGEM) calculation. The usual spectre-ml analysis then proceeds and the various voxel selections tested are then ranked by these metric values, with the best metric value indicating which voxel selection to use for the given application defined by the input spectra and choice of spectrum analysis and ranking metric. We note that no pyGEM code changes were required, indicating that spectre-ml end-users can relatively easily slot in their own analysis routines.
@@ II-B Uranium enrichment assay optimization
Multiple U3O8 standards measurements were performed with six US National Laboratory M400 detectors (Brookhaven, Idaho, Los Alamos, Oak Ridge, Pacific Northwest, and Sandia National Laboratories) [12] and provided to Lawrence Berkeley National Laboratory for analysis with spectre-ml. In this work we use the $30$-minute collimated M400 measurements of Standard Reference Material (SRM) 969 [13] samples with label U-235 enrichments of $1.94$ wt% and $4.46$ wt% (certified values of $(1.9420\pm 0.0014)$ wt% and $(4.4623\pm 0.0032)$ wt%). The latter is used as the known-enrichment sample for the pyGEM activity calibration while the former is used as the unknown sample to be assayed.
@@
The relative uncertainty in the U-235 enrichment is a crucial performance metric for U-235 NDA as it is directly used in computing the operator-inspector difference (OID) in terms of $z$-scores or “sigmas”, and used to compare uncertainties against the International Target Values (ITVs) for the GEM-based verification method [14]. A lower relative uncertainty on the measured U-235 enrichment increases the statistical power of the measurement to detect diversions from declared enrichments and thus potential treaty violations. Since the statistical component of the relative uncertainty decreases with the square root of measurement time, long measurement times may be required for high statistics; longer measurement times are however not always an option for inspectors performing verifications in the field. Optimizing for lower relative uncertainty can therefore improve the time-efficiency of IAEA NDA tasks.
@@
In the optimization examples of Section III, the spectre-ml optimization considers four different classes of algorithms for selecting voxel cluster combinations: machine learning (ML), heuristic, greedy, and random voxel selections. The ML parameter sweeps consist of $n_{\text{clus}}=2$66 clusters (Gaussian Mixture only), $n_{\text{comp}}=2$66 NMF components, and $\alpha_{W}=0.0$ (i.e., no NMF regularization). The heuristic algorithms are the equal-depth and edge-and-anode algorithms introduced in Ref. [5], with $n_{\text{clus}}=2$66 and $n_{\text{anode}}=15$, $n_{\text{edge}}=1$, respectively. The greedy detector algorithm is used, but not the greedy pixel, depth bin, or voxel algorithms due to memory constraints, and a single sample is run for each of the random pixel, depth bin, and voxel clusterers with $n_{\text{clus}}=2$66.
@@ III Results
Fig. 1 shows the results of the spectre-ml + pyGEM U-235 enrichment relative uncertainty optimization using the PNNL M400 uranium standards dataset. The enrichment relative uncertainty of $1.06\%$ from the bulk (unoptimized) spectrum is improved to $0.93\%$ (a relative improvement of $12\%$) when using $n_{\text{comp}}=3$ and retaining $1$ out of $3$ Gaussian Mixture clusters. This voxel mask reduces the $185.7$ keV net peak area from $8.22\times 10^{5}$ cps to $4.45\times 10^{5}$ cps (an efficiency reduction to $54\%$ of the unoptimized detector) and slightly degrades the peak full width at half maximum (FWHM) from $1.18\%$ to $1.20\%$, but improves the overall fit quality from $2.70$ to $1.93$. The top $5$ masks achieve similar final metrics of $0.93\%$$0.94\%$. Several of the clustering algorithms (notably the edge-and-anode) clusterer never achieve results better than the bulk value of $1.06\%$.
@@
Similar or better spectre-ml + pyGEM improvements in the U-235 enrichment relative uncertainty are observed across all six M400 detectors tested—see Table I. The metric relative improvements range from $12\%$ to $26\%$, with a mean of $20\%$. Five out of six detectors are best improved by a GaussianMixture clusterer with $2$ out of $3$ clusters retained, while the PNNL detector removes one additional cluster and the BNL detector sees best results from the random depth bin clusterer—see the best cluster masks in Fig. 2. The relative efficiencies after voxel cluster removal range from $54\%$ (PNNL) to $70\%$ (BNL), with the remaining four detectors more tightly grouped between $61\%$ and $68\%$. All optimizations produce very small FWHM degradations, on average changing from $1.15\%$ (bulk) to $1.21\%$ (best), for reasons that are not yet known. Finally, the pyGEM fit quality improves in all six optimizations, from an average of $2.81$ to $1.71$.
@@
Figure 3: Trends in pyGEM analysis metrics vs dwell time $t$, using the optimum long-dwell-computed mask from the INL (top row), SNL (middle row), and LANL (bottom row) detectors. Left column: U-235 enrichment (wt%). The optimized (orange) points are slightly offset along the $x$-axis from their corresponding bulk (blue) points for visual clarity. The gray band shows the SRM 969 certification of $1.9420\pm 0.0014$ wt% [13]. Right column: U-235 relative uncertainty (%).
@@ IV Discussion
The results of Section III retain a handful of limitations and present opportunities for future work. First, the spectre-ml parameter sweeps are much smaller than those used in Ref. [5], using only 2266 clusters and NMF components, only the Gaussian Mixture clusterer (among the ML options), and no NMF regularization. It is likely that the observed $20\%$ average relative improvement in the U-235 enrichment relative uncertainty metric would improve if the parameter sweeps were expanded. Currently, we are compute-limited by the worse performance of the underlying scikit-learn libraries on the Windows machine required for the pyGEM code, with even these smaller parameter sweeps taking >1{>}1 hr wall time on a $4$-core 11th Gen Intel Core i7-1185G7 processor. Performance is also degraded by the need to run both NMF and fitting routines on spectra of $O(1000)$ bins for the entire U-235 region rather than the $O(100)$ bins for a single photopeak. Better parallelization, either in the scikit-learn routines or over individual parameter combinations, could greatly improve the speed of the optimization.
We also note that the observed $20\%$ average relative improvement is smaller than the ${\sim}3\times$ improvement observed in the U-235 peak analysis of Ref. [5], as expected, due to the much lower systematic fit error arising from the more complex spectrum fit (which includes the $195$ keV peak) in pyGEM. This re-emphasizes the importance of carefully defining the performance metric to be used to avoid specification gaming.
@@
Finally, we note that the ease with which pyGEM was integrated with spectre-ml bodes extremely well for future integrations of other analysis codes. For instance, FRAM [15] could be integrated to test the ability of spectre-ml to improve the challenging analysis of multiple densely-packed photopeaks in Pu spectra.
@@ V Conclusions
We have integrated the pyGEM uranium enrichment analysis code with the spectre-ml spectroscopic optimization framework, and shown that the uranium enrichment relative uncertainty can be directly used as an optimization target. This spectre-ml + GEM workflow provides a $20\%$ relative improvement (averaged over six H3D M400 detectors) in the U-235 enrichment relative uncertainty in $30$-minute long measurements, and retains similar enrichment accuracy to the unoptimized pyGEM-only results. These reductions in enrichment relative uncertainty can lead to significant time savings, with the optimized results achieving the same overall relative uncertainty as the unoptimized $30$-minute measurements around $10\times$ faster.

### v2

@@ I Introduction
The H3D M400 gamma spectrometer [m400_spec_sheet] is being adopted by the International Atomic Energy Agency (IAEA) as its primary in-field uranium enrichment non-destructive assay (NDA) technology [iaea2022disp, lebrun2022next, dodane2023large]. The M400 features four CdZnTe (CZT) crystals each pixelated to an $11\times 11$ grid, and offers medium resolution gamma spectroscopy at room temperature in a compact form factor. In previous work, we showed that the spatial variations in detector performance [goodman2022energy] within the CZT crystal volumes could be exploited to optimize spectroscopic performance metrics [vavrek2025data]. In particular, a balance can be found between identifying and rejecting poorly-performing detector regions (e.g., the poor-resolution anode regions and/or poor-signal-to-background cathode regions as shown in Fig. 1, or pixels with high defect concentrations [bolotnikov2007cumulative]) and the associated loss of efficiency. Because of the large number of possible voxel combinations (${\sim}10^{7285}$), computing the globally optimum set of voxels to include is infeasible, and approximate, data-driven methods were developed. Several of the example optimizations therein focused on minimizing the relative uncertainty in a peak fit amplitude parameter as proofs of concept in lieu of more advanced spectral performance metrics.
Figure 1: Uranium gamma spectra as a function of depth in the PNNL M400 detector. Top: heatmap of all depth spectra. The white dotted lines indicate the depth bin selections used in the bottom panel. Bottom: bulk spectrum and spectra at three depth bin selections, each normalized to their intensity at $186$ keV (black dotted line) to better show their shapes. Depth bins near the anode degrade in resolution and can exhibit gain drifts, while depth bins near the cathode have increased backgrounds.
In this work, we have integrated our spectroscopic optimization software spectre-ml with GEM [berlizov2022gem], the actual software used by the IAEA to conduct uranium NDA measurements. We show that the relative uncertainty on the GEM enrichment calculation can directly be used as a spectre-ml optimization target, i.e., that spectre-ml can improve the statistical confidence of the IAEA’s in-field enrichment measurements without increasing measurement times, or conversely can achieve the same statistical confidence in shorter measurement times. Similarly, spectre-ml improvements could widen the pool of CZT crystals meeting the performance requirements necessary for in-field measurements, potentially driving down manufacturing costs. Such software-based optimization could reduce the need for and/or complement various hardware-based advances such as design stage optimizations [li2024research, §3.2], [huang2025machine] that do not account for individual crystal differences, digital pulse processing for improved energy resolution [petryk2025improved] [smith2024summary, §5.2], and the use of the more spatially-homogeneous CdZnTeSe [roy2019evaluation] instead of CZT.
@@ II-A Integration of the spectre-ml and pyGEM codes
The spectre-ml [vavrek2025data, vavrek2023spectral] code is a Python package for optimizing the performance of many-channel gamma spectrometers by intelligently rejecting data from poorly-performing detector channels (i.e., spatial regions). It uses non-negative matrix factorization (NMF) [lee1999learning, wang2012nonnegative] and unsupervised clustering algorithms to learn groups of voxels with similar performance, compute performance metrics across various voxel cluster combinations, and then sweep over hyperparameters such as clustering algorithm, number of clusters, number of NMF components, etc., to find the best overall set of voxels to include. Previous versions of spectre-ml were restricted to single Doniach [doniach1970many] or Gaussian peak fit workflows using the becquerel library for peak fits. In this work we have generalized the spectre-ml software to abstract out the dependence on becquerel and enable arbitrary user-defined spectrum analyses, in particular those based on GEM. The pyGEM [mcdonald2022gamma] code is a Python interface to the General Enrichment Measurements (GEM) code [berlizov2022gem] used by the IAEA for uranium enrichment NDA. It fits measured spectra in the ${\sim}120$$270$ keV region of interest using a set of known U-235 emission lines, forward scatter, and high-energy downscatter profiles, and has recently been updated to handle the asymmetric peak shapes of CZT. In contrast to our Doniach fit becquerel workflow, pyGEM uses a triple-Gaussian model for each peak in CZT—one primary peak, one low-energy tail, and one high-energy tail. The user provides pyGEM a spectrum from a sample of known U-235 enrichment, and it will then compute a linear calibration between the fit net count rate in the $185.7$ keV peak and the U-235 atom abundance that can be used to compute the enrichment from an unknown sample spectrum. It is also possible to incorporate correction factors for certain changes between the calibration and sample spectra such as changes in the sample matrix (e.g., U metal vs. U3O8), Al and steel wall thicknesses, etc. The pyGEM enrichment relative uncertainty calculation incorporates both statistical and systematic (i.e., fit error) components, and thus the spectre-ml optimization can be viewed as trading off statistical vs. systematic uncertainties as voxel clusters with higher fit error are removed.
@@
To integrate the two codes, some spectre-ml updates were required to handle both the becquerel and pyGEM workflows. As shown in the pseudocode in the Appendix, spectre-ml now provides the abstract base classes SpectrumAnalyzer and RankingMetric. The user must provide concrete subclasses implementing SpectrumAnalyzer.analyze and RankingMetric.calc, where the latter method takes a spectrum analyzed (e.g., fit) by the former and extracts a single float metric such as the relative uncertainty in a fit parameter (becquerel) or in an enrichment (pyGEM) calculation. The usual spectre-ml analysis then proceeds and the various voxel selections tested are then ranked by these metric values, with the best metric value indicating which voxel selection to use for the given application defined by the input spectra and choice of spectrum analysis and ranking metric. An overview of the pipeline is given in Fig. 2. We note that no pyGEM code changes were required, indicating that spectre-ml end-users can relatively easily slot in their own analysis routines.
Figure 2: Overview of the spectre-ml + pyGEM pipeline, adapted from Fig. 2 of Ref. [vavrek2025data].
@@ II-B Uranium enrichment assay optimization
Multiple U3O8 standards measurements were performed with six US National Laboratory M400 detectors (Brookhaven, Idaho, Los Alamos, Oak Ridge, Pacific Northwest, and Sandia National Laboratories) [smith2024summary] and provided to Lawrence Berkeley National Laboratory for analysis with spectre-ml. In this work we use the $30$-minute collimated M400 measurements of Standard Reference Material (SRM) 969 [srm969] samples with label U-235 enrichments of $1.94$ wt% and $4.46$ wt% (certified values of $(1.9420\pm 0.0014)$ wt% and $(4.4623\pm 0.0032)$ wt%). The latter is used as the known-enrichment sample for the pyGEM activity calibration while the former is used as the unknown sample to be assayed.
@@
The relative uncertainty in the U-235 enrichment is a crucial performance metric for U-235 NDA as it is directly used in computing the operator-inspector difference (OID) in terms of $z$-scores or “sigmas”, and used to compare uncertainties against the International Target Values (ITVs) for the GEM-based verification method [iaea2022itvs]. A lower relative uncertainty on the measured U-235 enrichment increases the statistical power of the measurement to detect diversions from declared enrichments and thus potential treaty violations. Since the statistical component of the relative uncertainty decreases with the square root of measurement time, long measurement times may be required for high statistics; longer measurement times are however not always an option for inspectors performing verifications in the field. Optimizing for lower relative uncertainty can therefore improve the time-efficiency of IAEA NDA tasks.
It is also useful to quantify improvements in the pyGEM fit quality after voxel clusters are removed. A pyGEM analysis typically uses the “Nqfit” metric, an “intensity-normalized” version of the standard reduced chi-squared $\chi_{\nu}^{2}$, given by [iaea2020physical]
${\tt Nqfit}-1=\left(\chi_{\nu}^{2}-1\right)\frac{100\times 10^{3}}{{\tt sum}+50\times 10^{3}}.$ (1)
Here sum is the total counts of the U-235 fit component, such that Nqfit is essentially a $\chi_{\nu}^{2}$ scaled to an assumed “normal” counting statistics of $50\times 10^{3}$ counts.
In the optimization examples of Section III, the spectre-ml optimization considers four different classes of algorithms for creating voxel clusters (see also Ref. [vavrek2025data]):
  1. 1.
    machine learning (ML): standard machine learning clustering algorithms, e.g., Gaussian Mixture, Agglomerative, or $k$-means clustering, using the NMF spectral decomposition weights as inputs;
  2. 2.
    heuristic:
    1. (a)
      equal-depth: segments the detector into $n_{\text{clus}}$ equally-sized regions in depth;
    2. (b)
      edge-and-anode: segments the detector into three regions: a depth region of $n_{\text{anode}}$ depth bins, an edge region of the $n_{\text{edge}}$ outer pixels of each of the four detector modules (not including the anode depth bins), and the rest of the detector volume;
  3. 3.
    greedy: computes the metric directly on spectra within every distinct spatial element (voxel, pixel, depth bin, or crystal), ranks each element by its individual metric, then successively clusters elements together in that order;
  4. 4.
    random: assigns $n_{\text{clus}}$ random cluster labels at the voxel, pixel, depth bin, or crystal level.
@@
The ML parameter sweeps consist of $n_{\text{clus}}=2$66 clusters (Gaussian Mixture only), $n_{\text{comp}}=2$66 NMF components, and $\alpha_{W}=0.0$ (i.e., no NMF regularization). The equal-depth and edge-and-anode algorithms use $n_{\text{clus}}=2$66 and a fixed $(n_{\text{edge}},n_{\text{anode}})=(1,15)$, respectively. The greedy detector algorithm is used, but not the greedy pixel, depth bin, or voxel algorithms due to memory constraints, and a single sample is run for each of the random pixel, depth bin, and voxel clusterers with $n_{\text{clus}}=2$66.
@@ III Results
Fig. 3 shows the results of the spectre-ml + pyGEM U-235 enrichment relative uncertainty optimization using the PNNL M400 uranium standards dataset. The enrichment relative uncertainty of $1.06\%$ from the bulk (unoptimized) spectrum is improved to $0.93\%$ (a relative improvement of $12\%$) when using $n_{\text{comp}}=3$ and retaining $1$ out of $3$ Gaussian Mixture clusters. This voxel mask reduces the $185.7$ keV net peak area from $8.22\times 10^{5}$ cps to $4.45\times 10^{5}$ cps (an efficiency reduction to $54\%$ of the unoptimized detector) and slightly degrades the peak full width at half maximum (FWHM) from $1.18\%$ to $1.20\%$, but improves the overall fit quality Nqfit from $2.70$ to $1.93$ by reducing fit residuals in both continuum and photopeak regions. The top $5$ masks achieve similar final metrics of $0.93\%$$0.94\%$. The Gaussian Mixture clusterer generally outperforms the heuristic clustering algorithms, since it can learn and better represent trends in a particular dataset. The edge-and-anode and random pixel and voxel clusterers achieve results worse than the bulk value of $1.06\%$, due in part to their inability to learn from the data, but also their more limited parameter sweeps.
@@
Similar or better spectre-ml + pyGEM improvements in the U-235 enrichment relative uncertainty are observed across all six M400 detectors tested—see Table I. The metric relative improvements range from $12\%$ to $26\%$, with a mean of $20\%$. Four out of six detectors are best improved by a GaussianMixture clusterer with $2$ out of $3$ clusters retained, while the PNNL detector removes one additional cluster and the BNL detector sees best results from the random depth bin clusterer—see the best cluster masks in Fig. 4. These six masks, though found for different detectors via different model parameters, all remove largely depth-based clusters, mostly near the cathode and/or anode. CZT resolution degradation near the anode is well-known [li1999spatial], and worse signal-to-background ratios are observed near the cathode in Fig. 1. See also the discussion on the similarity of the six masks in Section IV. The relative efficiencies after voxel cluster removal range from $54\%$ (PNNL) to $70\%$ (BNL), with the remaining four detectors more tightly grouped between $61\%$ and $68\%$. All optimizations produce very small FWHM degradations, on average changing from $1.15\%$ (bulk) to $1.21\%$ (best), for reasons that are not yet known. Finally, the pyGEM fit quality improves in all six optimizations, from an average Nqfit of $2.81$ to $1.71$.
@@
Figure 5: Trends in pyGEM analysis metrics vs dwell time $t$, using the optimum long-dwell-computed mask from the INL (top row), SNL (middle row), and LANL (bottom row) detectors. Left column: U-235 enrichment (wt%). The optimized (orange) points are slightly offset along the $x$-axis from their corresponding bulk (blue) points for visual clarity. The gray band shows the SRM 969 certification of $1.9420\pm 0.0014$ wt% [srm969]. Right column: U-235 relative uncertainty (%).
TABLE II: Summary of time efficiency improvements
@@ IV Discussion
The results of Section III demonstrate that spectre-ml can improve uranium enrichment assay uncertainties found via the pyGEM code, thereby offering faster measurement times for in-field safeguards inspections. Here we discuss a handful of remaining limitations, and present opportunities for future work.
First, the spectre-ml parameter sweeps are much smaller than those used in Ref. [vavrek2025data], using only 2266 clusters and NMF components, only the Gaussian Mixture clusterer (among the ML options), and no NMF regularization. It is likely that the observed $20\%$ average relative improvement in the U-235 enrichment relative uncertainty metric would improve if the parameter sweeps were expanded. Currently, we are compute-limited by the worse performance of the underlying scikit-learn libraries on the Windows machine required for the pyGEM code, with even these smaller parameter sweeps taking >1{>}1 hr wall time on a $4$-core 11th Gen Intel Core i7-1185G7 processor. Performance is also degraded by some difficult-to-avoid array copies in the pyGEM pipeline, as well as the need to run both NMF and fitting routines on spectra of $O(1000)$ bins for the entire U-235 region rather than the $O(100)$ bins for a single photopeak. Better parallelization, either in the scikit-learn routines or over individual parameter combinations, could greatly improve the speed of the optimization.
We also note that the observed $20\%$ average relative improvement is smaller than the ${\sim}3\times$ improvement observed in the U-235 peak analysis of Ref. [vavrek2025data], as expected, due to the much lower systematic fit error arising from the more complex spectrum fit (which includes the $195$ keV peak) in pyGEM. This re-emphasizes the importance of carefully defining the performance metric to be used to avoid specification gaming.
@@
Finally, we note that the ease with which pyGEM was integrated with spectre-ml bodes extremely well for future integrations of other analysis codes and/or other detector arrays. For instance, FRAM [sampson1989fram] could be integrated to test the ability of spectre-ml to improve the challenging analysis of multiple densely-packed photopeaks in Pu spectra.
@@ V Conclusions
We have integrated the pyGEM uranium enrichment analysis code with the spectre-ml spectroscopic optimization framework, and shown that the uranium enrichment relative uncertainty can be directly used as an optimization target. This spectre-ml + GEM workflow provides a $20\%$ relative improvement (averaged over six H3D M400 detectors) in the U-235 enrichment relative uncertainty in $30$-minute long measurements, and retains similar enrichment accuracy to the unoptimized pyGEM-only results. These reductions in enrichment relative uncertainty can lead to significant time savings for in-field nuclear safeguards inspections, with the optimized results achieving the same overall relative uncertainty as the unoptimized $30$-minute measurements around $10\times$ faster. Future work could integrate other safeguards NDA software such as FRAM with spectre-ml for improved Pu spectroscopy, and investigate whether a single common optimization result can be applied across multiple detectors to provide improved spectroscopy with minimal operational cost.